OpenStudy (anonymous):

What is the equation of a line with a point (-3, 0) and no slope? x + 3 = 0 y + 3 = 0 x - 3 = 0 y - 3 = 0

OpenStudy (anonymous):

in this case the line is parallel to x axis. y=mx+c y=-3 y+3=0 answer

OpenStudy (redprince):

\[y+3=0\]